A rock is dropped from rest off of a high cliff on another planet, planet X. There is no atmosphere, and thus no air resistance, on planet X. At the end of 1 second, the rock is moving at a speed of 6 m/s. What is the rock's average speed during the entire first second [from t=0 until t=1 sec]?

6 m/s

0

10 m/s

5 m/s

3 m/s
Answer #1

Initial velocity, u = 0
Final velocity, v = 6 m/s
Time taken, t = 1 s
Acceleration, a = (v - u) / t
= (6 - 0) / 1
= 6 m/s2

Distance travelled in 1 s, s = u * t + 1/2 a * t2
= 0 + 0.5 * 6 * 12
= 3 m

Average velocity = s/t
= 3 m / 1 s
= 3 m/s